3x-2y=-9
X-5y=-16
What is the y-coordinate of the point of the intersection of these two lines?

Answer by stanbon(75887)

3x-2y=-9
X-5y=-16
What is the y-coordinate of the point of the intersection of these two lines?
---
Solve each eq. for "x":
x = (2/3)y -3
x = 5y-16
Substitute for "x" and solve for "y"
(2/3)y - 3 = 5y-16
Multiply thru by 3 to get:
2y - 9 = 15y - 48
13y = +39
y = 3
========================
Cheers,
Stan H.